Week of 170522

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:

  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Andrea (MW Officer + data mgmt), Ben (computing), David C (databases), Julia (WLCG), Kate (WLCG), Maarten (SCOD + ALICE + GGUS), Marian (networks + ETF), Paul (storage), Vincent (security), Xavi (storage)
  • remote: David B (IN2P3), David M (FNAL), Dimitri (KIT), Gareth (RAL), Kyle (OSG), Luca (CNAF), Onno (NLT1), Stefano (CMS), Victor (JINR), Vincenzo (EGI), Xin (BNL), Zoltan (LHCb)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Activities:
      • MC simulation dominating between ATLAS activities
      • End of reprocessing campaign with minority trigger stream processing
      • Preparation for data-taking
    • Issues:
      • Overlay tasks caused trouble to Squids and Frontier servers again during the week-end but failover system worked OK.

  • CMS reports ( raw view) -
    • I may be on the road at the time of the meeting, plan to connect via Vidyo client on my phone. I apologize in advance in case that will not work
    • Load on computing grid still a bit low due to lack of requests.
    • Only one major problem this week.
      • ALARM ticket GGUS:128348 Argus stopped working for CMS on Monday, May 15, afternoon with PEP calls timing out. Made SAM tests fail at CERN site and stopped analysis job submission for all of CMS. CMS identified problem in the evening, opened alarm at 11 pm, expert got to it next morning at 7:30 am and fixed it. We do not know yet what the problem was, nor why nobody else was apparently affected. Argus was green in SLS.
        • Maarten: also ALICE were very much affected
        • Ben: the trouble arose when a new VO was added incorrectly,
          which should not have caused problems for existing VOs;
          we have followed up with the devs
        • Stefano: there ought to be an SLS probe that imitates what clients do
        • Ben: yes, it appears that none of the current probes exposed the problem
      • Reminder: CMS Remote Analysis Builder (CRAB) relies on CERN GSI authentication to map every DN into a different user when submitting to HTCondor.

  • ALICE -
    • continued high activity

  • LHCb reports ( raw view) -
    • Activity
      • User analysis and Monte Carlo simulation
      • New validation of Stripping v24 has been started.
    • Site Issues
      • T1:
        • RAL: disk server failures during the weekend
    • Zoltan: we are looking into an issue with the latest gfal-utils vs. a json-c dependency

Sites / Services round table:

  • ASGC:
  • BNL: Ran a tape staging stress test, hit a bug in the new tape interface of dCache3, contacted dCache developers.
  • CNAF:
    • On Friday we had a problem with the main mail relay, it was shut down by accident.
  • EGI: NTR
  • FNAL: NTR
  • IN2P3: NTR
  • JINR: We lost main link to Moscow around 11:00 MSK 20-05-2017.Unfortunately backup link is not able to support JINR full speed connectivity at present. The problem is being worked on.
  • KISTI:
  • KIT: NTR
  • NDGF: Cooling failure at UCPH during weekend; Some computation nodes died. Downtime for HPC2N computation nodes all working hours Tuesday (only ATLAS affected). We should still be able to fulfill pledge. (Christian: I have connection issues with vidyo)
  • NL-T1: Scheduled dCache upgrade to 2.16 for the week of June 6: https://goc.egi.eu/portal/index.php?Page_Type=Downtime&id=23111 and https://goc.egi.eu/portal/index.php?Page_Type=Downtime&id=23112
  • NRC-KI:
  • OSG:
    • Kyle: GGUS:127855 about intermittent SAM tests should be routed to the SAM experts
      • Marian: will look into it
  • PIC:
  • RAL: Updates of Castor to version 2.1.16 for Atlas on Tuesday (tomorrow) and CMS Thursday (25th).
    • Gareth: last week our new UPS was tested OK
  • TRIUMF:NTR

  • CERN computing services:
  • CERN storage services:
    • EOS Alice crashed at lunch time today. It restarted automatically. We are looking for what exactly happened.
  • CERN databases: Nothing to report
  • GGUS: NTR
  • Monitoring:
    • Final reports for the April availability sent around.
  • MW Officer:
    • dCache 2.13.x end of support from 1st of June. 27 instances running this version ( Triumf, NL_T1, KIT, IN2P3, FNAL as T1 ).
  • Networks: NTR
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r21 < r20 < r19 < r18 < r17 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r21 - 2017-05-23 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback